An investment project provides cash inflows of $675 per year for 8 years. What is the project payback period if the initial cost is $1,700?
Answer #1

Payback period of the project can be calculated-

Outflows at the beginning= 1700

Payment in first two installment= (675*2)=1350

Payback period = [2 years+(1700-1350)/675]

= 2+(350/675)

= 2.5185